# Learning from examples: AI assistance can enhance rather than hinder skill development

### v1

@@ Abstract
It is widely believed that outsourcing cognitive work to AI boosts immediate productivity at the expense of long-term human capital development. An opposing possibility is that AI tools can support skill development by providing just-in-time, high-quality, personalized examples. This work explores whether using an AI writing tool undermines or supports performance on later unaided writing. In Study 1, forecasters predicted that practicing writing cover letters with an AI tool would impair learning compared to practicing alone. However, in Study 2, participants randomly assigned to practice writing with AI improved more on a subsequent writing test than those assigned to practice without AI (d = 0.40***)—despite exerting less effort, whether measured by time on task, keystrokes, or subjective ratings. In Study 3, participants who had practiced writing with AI again outperformed those who practiced without AI (d = 0.31***). Consistent with the positive impact of exposure to high-quality examples, these participants performed just as well as those who viewed–but could not edit–an AI-generated cover letter (d = 0.03, ns). In both Studies 2 and 3, the benefits of practicing with AI persisted in a one-day follow-up writing test. Collectively, these findings constitute an existence proof that, contrary to participants’ intuition, using AI tools can improve, rather than undermine, learning.
@@
Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6). For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI will “lead to humans becoming less intelligent” (7). In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (8).” When this ban was lifted three months later, it was not because of the potential of AI to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (9).” The sentiment behind the initial ban aligns with teacher perceptions: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (10).
@@
Third, technological tools reduce the need for the learner to be cognitively engaged with the task at hand. For instance, knowing that we will be able to search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (14). And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (15, 16). To the extent that tools powered by generative AI instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning. It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
There is a plausible, albeit less obvious, alternate hypothesis. Using AI tools to do our work alongside us could help us develop our own skills. In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp and apply. Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (17, 18). Compared to textbooks, conventional computer tutoring programs (e.g., (19)), and even human teachers, today’s AI tools are unprecedented in their ability to deliver hyper-personalized high-quality examples on command. Thus, AI tools may improve skill development if the benefits of exposure to just-in-time examples tailored to learners’ needs outweigh the costs of diminished engagement.
There is little research on the effect of AI tools on subsequent skill development, as opposed to productivity during use. In working papers, results have been mixed. Some studies have found that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (20, 21), while others have shown null or even negative effects (22, 23, 21). Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions as is typical in real-world use. Further, they focus exclusively on mathematics and computer programming.
@@
To differentiate the effects of AI use on learning versus performance, we developed a paradigm in which all participants were given a baseline writing test (i.e., revising a poorly written cover letter). They were then introduced to evidence-based strategies for professional writing, with descriptions and examples for each one (25, 26). Next, participants were randomly assigned to one of three conditions: practicing rewriting a different cover letter with access to an AI tool based on the same writing principles that users had just learned (See Figure 1), practicing without this tool, or a no-practice control group. To assess gains in writing skill, participants completed an incentivized test in which they rewrote yet another cover letter without access to AI, with a cash bonus guaranteed for submissions ranked in the top 10 percent. To assess the persistence of skill improvement, all participants were invited to complete a similar incentivized test of writing skill one day later. See Figure 2. We used GPT4o to rate each cover letter for each of the five writing principles introduced in this experiment. We averaged these ratings to produce summary scores of writing skill and, in a random subsample (n = 30), validated these scores using trained human raters (r = .83, p < .001). Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview. The relative likelihood of a cover letter securing an interview correlated positively with AI ratings of writing skill (rs = .29 and .28, ps < .001, for Study 2 and 3, respectively).
@@
The learning advantage of having practiced with AI was evident in the test phase: Consistent with our pre-registered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001). See Figure S4. Likewise, cover letters written by participants who had practiced with AI were more likely to secure a hypothetical job interview than cover letters by participants who had practiced without AI (.54 vs. .50, p = .002) or had not practiced at all (.54 vs. .47, p < .001). See Figure 5.
@@
To gain insight into what might be driving the benefit of practicing with AI, in Study 3 (N = 2003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition. In this condition, participants were shown an AI-generated writing example that they could not edit. To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
@@
As in Study 2, participants given access to the AI writing tool dramatically outperformed participants who did not get access to it, both while using it during the practice phase (d = 1.01, p <.001), and during the no-AI test phase (d = .28, p < .001). Their test-phase cover letters were also relatively more likely to secure them hypothetical job interviews (.51 vs. .47, p = .007).
Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .31, p = < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs. .47, p = .007). Notably, they improved as much as participants who had practiced with the AI tool (and could edit its output, d = .03, p = .883), and were offered hypothetical interviews at similar rates as them (.51 vs. .52, p = .561). See Figures S6 and 7.
@@
During the test phase, participants who had seen the AI example worked for an additional 55 seconds more than participants who had practiced without AI (d = 0.22, p < .001). They also logged 30% more keystrokes than participants who had practiced with AI (d = 1.86 p <.001) and 48% more than those who practiced without AI (d = 2.39, p < .001). Across conditions, all participants self-reported similar levels of subjective effort (ds < .08, ps > .05).
@@
When we recontacted a subsample of (n = 800) participants one day later, the majority responded (n = 618, 77%); the attrition rates ranged from 17% to 24% and did not differ by condition ($\sf\chi^{2}$ (2) = 4.56, p = .102). The effect remained robust after 24 hours. Participants who had practiced with the AI-tool (d = .32, p = .005) and participants who had simply seen an AI example (d = .39, p < .001), both continued to outperform those who had practiced without the tool. Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .07, p = .745). See Figure 6 and Section C5 of Supplementary Information for details.
@@ Discussion
Contrary to the expectations of people shown our paradigm (Study 1), access to the AI writing tool led to improvements in writing skill (unassisted by AI) that persisted in a one-day follow-up test (Study 2). These learning gains cannot be attributed to greater effort, as participants who practiced with AI expended less effort during practice than those who practiced alone. Instead, these gains may be explained by exposure to a high-quality, just-in-time personalized example: participants who merely viewed an AI-generated example cover letter (without editing it) improved their writing skill just as much as those who given the option to practice editing the cover letter (Study 3).
While not obvious to forecasters (Study 1), teachers (10), and the general population (7), the benefits of seeing AI-generated examples are consistent with prior research on how people learn. In addition to the literature on worked examples mentioned earlier, research has shown that humans are especially adept at observing, imitating, and learning from others (27, 28, 29). Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (30).
@@
We want to highlight three promising directions for future research. First, we focused on the modal use case for AI tools in the workplace—professional writing (24). It is unclear whether learning from AI examples occurs in other domains. Indeed, in mathematics and computer programming, AI tools do not always support learning (22, 23). Indeed, writing may be particularly suited to learning by example. At a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., shorter sentences and the strategic use of line breaks and bold-face formatting). In other domains, merely observing a solution may be less informative. For instance, the final answer to a math problem does not reveal the procedure that produced it.
@@
Third, in our experimental paradigm, participants who interacted with the AI tool did so only once. It is common, however, to use AI tools repeatedly. When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time? Future research, ideally in field settings, is needed to establish the long-term benefits and costs of relying on AI tools.
@@ Conclusion
Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development. Although it reduced the effort users invested in practicing, the AI writing tool nevertheless accelerated skill development. Id accomplished this by providing high-quality, just-in-time, personalized examples of excellent writing. The underappreciated efficacy of timely and tailored examples has practical implications for the design of AI tools. Many AI tools designed to support learning are explicitly programmed not to “give away” answers. It may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
@@ Methods
As per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality. To do this, we independently rated each cover letter and each writing principle. Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (32, 33). See our prompts in Table S1. We then took the unweighted average of these 5 scores as our main dependent variable. See disaggregated analyses by each writing principle and additional outcomes on Table S6 and S16.

### v2

# Learning not cheating: AI assistance can enhance rather than hinder skill development

@@ Abstract
It is widely believed that outsourcing cognitive work to AI boosts immediate productivity at the expense of long-term human capital development. An overlooked possibility is that AI tools can support skill development by providing just-in-time, high-quality, personalized examples. In this investigation, lay forecasters predicted that practicing writing cover letters with an AI tool would impair learning compared to practicing writing letters without the tool. However, in a highly-powered pre-registered experiment, participants randomly assigned to practice writing with AI improved more on a writing test one day later compared to writers assigned to practice without AI. Notably, writers given access to the AI tool improved more despite exerting less effort, whether measured by time on task, keystrokes, or subjective ratings. We replicated and extended these results in a second pre-registered experiment, showing that writers given access to the AI tool again outperformed those who practiced on their own—but performed no better than writers merely shown an AI-generated cover letter that they could not edit. Collectively, these findings constitute an existence proof that by providing personalized examples of high-quality work, AI tools can improve, rather than undermine, learning.
keywords:
Human Capital Development | Generative AI | Skill Acquisition
@@
Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6, 7). For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI will “lead to humans becoming less intelligent” (8). In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of the potential of AI to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” The sentiment behind the initial ban aligns with teacher perceptions: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
@@
Third, technological tools reduce the need for the learner to be cognitively engaged with the task at hand. For instance, knowing that we will be able to search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15). And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (16, 17). To the extent that tools powered by generative AI instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning. It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
There is a plausible, albeit less obvious, alternate hypothesis. Using AI tools to help do our work could help us develop our own skills. In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp and apply. Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (18, 19). Compared to textbooks, conventional computer tutoring programs (e.g., (20)), and even human teachers, today’s AI tools are unprecedented in their ability to deliver hyper-personalized high-quality examples on command. Thus, AI tools may improve skill development if the benefits of exposure to just-in-time examples tailored to learners’ needs outweigh the costs of diminished engagement.
There is little research on how AI tools influence skill development. In working papers, results have been mixed. Some studies have found that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22). Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions as is typical in real-world use. Further, they focus exclusively on mathematics and computer programming.
@@
To differentiate the effects of AI use on learning versus performance, we developed a paradigm in which all participants were given a baseline writing test (i.e., revising a poorly written cover letter). They were then introduced to evidence-based strategies for professional writing, with descriptions and examples for each one (26, 27). Next, participants were randomly assigned to one of three conditions: practicing rewriting a different cover letter with access to an AI tool based on the same writing principles that users had just learned (See Figure 1), practicing without this tool, or a no-practice control group. To assess gains in writing skill, participants completed an incentivized test in which they rewrote yet another cover letter without access to AI, with a cash bonus guaranteed for submissions ranked in the top 10 percent. To assess the persistence of skill improvement, all participants were invited to complete a similar incentivized test of writing skill one day later. See Figure 2. We used GPT4o to rate each cover letter for each of the five writing principles introduced in this experiment. We averaged these ratings to produce summary scores of writing skill and, in a random subsample (n = 30), validated these scores using trained human raters (r = .83, p < .001). Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview. The relative likelihood of a cover letter securing an interview correlated positively with AI ratings of writing skill (rs = .29 and .28, ps < .001, for Study 2 and 3, respectively).
@@
The learning advantage of having practiced with AI was evident in the test phase: Consistent with our pre-registered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001). See Figure S4. Likewise, cover letters written by participants who had practiced with AI were more likely to secure a hypothetical job interview than cover letters by participants who had practiced without AI (.54 vs. .50, p = .002) or had not practiced at all (.54 vs. .47, p < .001). See Figure 5. Notably, the effect of practicing with AI was not limited to stronger or weaker writers—participants across all skill levels benefited equally from practicing with AI.
@@
To gain insight into what might be driving the benefit of practicing with AI, in Study 3 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition. In this condition, participants were shown an AI-generated writing example that they could not edit. To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
@@
As in Study 2, participants given access to the AI writing tool dramatically outperformed participants who did not get access to it, both while using it during the practice phase (d = 1.22, p <.001), and during the no-AI test phase (d = .34, p < .001). Their test-phase cover letters were also relatively more likely to secure them hypothetical job interviews (.51 vs. .47, p = .008).
Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p = < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs. .47, p = .007). Notably, they improved as much as participants who had practiced with the AI tool (and could edit its output, d = .03, p = .883), and were offered hypothetical interviews at similar rates as them (.51 vs. .52, p = .561). See Figures S6 and 7. As in Study 2, gains from practicing with AI and seeing AI examples were comparable across all levels of baseline writing skill.
@@
During the test phase, participants who had seen the AI example worked for an additional 56 seconds more than participants who had practiced without AI (d = 0.22, p < .001). They also logged 30% more keystrokes than participants who had practiced with AI (d = 1.86 p <.001) and 60% more than those who practiced without AI (d = 2.39, p < .001). Across conditions, all participants self-reported similar levels of subjective effort (ds < .08, ps > .05).
@@
When we recontacted a subsample of (n = 800) participants one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ($\sf\chi^{2}$ (2) = 4.56, p = .102). The effect remained robust after 24 hours. Participants who had practiced with the AI-tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool. Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830). See Figure 6 and Section C5 of Supplementary Information for details.
@@ Discussion
Contrary to the expectations of lay forecasters (Study 1), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 3). Specifically, participants who had practiced with AI wrote (without AI assistance) cover letters that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later. Learning gains were not the result of greater effort; in fact, participants who’d practiced with AI expended less effort during practice than those who’d practiced alone. Instead, these gains can be explained by exposure to a high-quality, just-in-time personalized example: participants who merely viewed an AI-generated example cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 3).
While not obvious to forecasters, teachers (11), and the general population (8), the benefits of seeing AI-generated examples are consistent with prior research on how people learn. In addition to the literature on worked examples mentioned earlier, research has shown that humans are especially adept at observing, imitating, and learning from others (28, 29, 30). Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (31).
@@
Three promising directions for future research are worth highlighting. First, it remains to be seen whether learning from AI examples occurs in domains other than writing, currently the modal use case for AI tools in the workplace (25) . Writing may be particularly suited to learning by example because at a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting). In other domains, merely observing a solution may be less informative. For instance, the final answer to a math or computer programming problem does not instantly reveal the procedure that produced it. Studies of AI tutors in these domains have, perhaps not surprisingly, found null or negative effects (23, 24).
@@
Third, in our experimental paradigm, participants who interacted with the AI tool did so only once. It is common, however, to use AI tools repeatedly. When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time? Consider, for instance, the game of Go. The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (32). Future research, ideally in field settings, is needed to establish the long-term benefits and costs of relying on AI tools.
@@ Conclusion
Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development. Although it reduced the effort users invested in practicing, the AI writing tool nevertheless accelerated skill development. It accomplished this by providing high-quality, just-in-time, personalized examples of excellent writing. The underappreciated efficacy of timely and tailored examples has practical implications for the design of AI tools. Many AI tools designed to support learning are explicitly programmed not to “give away” answers. It may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
@@ Methods
As per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality. To do this, we independently rated each cover letter and each writing principle. Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (34, 35). See our prompts in Table S1. We then took the unweighted average of these 5 scores as our main dependent variable. See disaggregated analyses by each writing principle and additional outcomes on Table S6 and S16.